OpenStudy (anonymous):

y=5x^2+7x-5 Use the quadratic formula to find any x-intercept on the graph of the equation.

OpenStudy (campbell_st):

the quadratic formula is use for quadratics \[ax^2 + bx + c = 0\] \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] in your question a the x-intecepts are where y = 0 so you would have a = 5, b = 7 and c = -5 just substitute and evaluate.

OpenStudy (anonymous):

how would i do that

OpenStudy (anonymous):

can you help me with that

OpenStudy (precal):



OpenStudy (precal):

do order of operation

OpenStudy (anonymous):

is that the x intercept

OpenStudy (precal):

yes

OpenStudy (precal):

what number is under the square root sign